A loan of P5,000 is made for a period of 15 months, at a simple interest rate of 15%, what future amount is due at the end of the loan period?


1
Expert's answer
2022-02-28T08:31:45-0500

"I = P \times r \times t"

t= "\frac{15}{12}= 1.25"

"I = 5000\times 0.15 \times 1.25"

= "937.5"

Total Amount to be paid= Interest+ Principle amount

="937.5+ 5000= P5937.50"